OpenStudy (anonymous):

Differential calculus mixing problem: There is a dam with a volume of 49 cubic metres. Water flows out of this into one of the creeks. Salt water with salt content of 30kg/m3 flowed into the dam at 0.68m3/s and mixed with the fresh water. The mixed water flowed out at the same rate. Flag tail in the dam can tolerate salinity up to 14kg/m3. How long after the flow starts can the fish survive in the dam?

OpenStudy (anonymous):

Do you know the formula for mixtures?

OpenStudy (anonymous):

I can show you my working so far.

OpenStudy (anonymous):

\[\frac{ dQ }{ dt }= Rate In-RateOut\]

OpenStudy (anonymous):

Rate In = (Concentration in)(Flow Rate In) Rate Units = lb/min Concentration Units = lb/gal Flow Units = gal/min Other units may be used if consistent throughout.

OpenStudy (anonymous):

I know the formula, and i have attempted the question but the answer i got was 0.045 seconds which must be wrong.

OpenStudy (anonymous):

input: 20.4Kg/s Output: 0.01387755x

OpenStudy (nincompoop):

don't you need to solve for t?

OpenStudy (anonymous):

i cant be sure i have solved the differential equation correctly. as : \[\frac{ dx }{ dt}=20.4 - 0.01387755x\]

OpenStudy (anonymous):

How did you get your units output and input above?

OpenStudy (anonymous):

concentration x flow rate

OpenStudy (anonymous):

How would i go about doing that, don't i have to have a completed model to solve for t.

OpenStudy (anonymous):

Lets take it step by step,\[Rate In= (30kg/m^3)(0.63m^3/s) = 18.9\frac{ kg }{ s }\]

OpenStudy (anonymous):

20.4

OpenStudy (anonymous):

0.68, = 20.4Kg/s

OpenStudy (anonymous):

My apology I just noticed.

OpenStudy (anonymous):

All good

OpenStudy (anonymous):

if this is correct so far,do i just solve the differential equation and find t?

OpenStudy (anonymous):

\[RateOut = (\frac{ y(t)kg }{ 49m^3 })(0.68kg/s)\]

OpenStudy (nincompoop):

https://www.youtube.com/watch?v=bqVzRBf_NGM

OpenStudy (anonymous):

Not sure as of yet since the 14kg/m^3 may be used to determine the value for C that we find to actually solve for the value of t

OpenStudy (anonymous):

I cannot access youtube, i am using a school computer atm. thanks though.

OpenStudy (anonymous):

Another thing, is that the fish can only tolerate a salinity of 14Kg/m^3. Does this not mean that the amount of salt in the dam would actually be 686Kg as this creates a concentration in the pond equal to 14Kg/m^3? 14 x 49 = 686

OpenStudy (anonymous):

It actually may be that we use 14kg/m^3 for the concentration out since that is what we are looking for at a time.

OpenStudy (anonymous):

Then solving the equation you will end up with a 2 variables T and C. To solve for C you will use the fresh water initially or Y(0) = 49m^3 so the T will go away and you will find a value for C and from there all you would need to do is algebraically solve for t.

OpenStudy (anonymous):

Ok, i'll give it s go.

OpenStudy (anonymous):

Ninc, do you know if this is correct?

OpenStudy (anonymous):

Doesn't look correct no but even so you did not use the intial value Y(o) = 49m^3 so that C will be 49m^3

OpenStudy (anonymous):

In most of the examples i found similar the final model contained e

OpenStudy (anonymous):

Correct so have I, that is why I am not seeing how this problem is going so far and the units are not helping me either

OpenStudy (anonymous):

http://www4.ncsu.edu/~msolufse/MA432ODEs.pdf

OpenStudy (anonymous):

On page 10 of this pdf there is a very similar question, you may be able to follow it better than i did.

OpenStudy (anonymous):

We initially followed that.